# Lightweight Dataset Pruning without Full Training via Example Difficulty and Prediction Uncertainty

### v2

@@ A.2 Detailed Experimental Settings
Here we clarify the technical details of our work. For training the model on the full dataset and the selected subset, all parameters are used identically except for batch sizes. For CIFAR-10/100, we train ResNet-18 for 200 epochs with a batch size of 128, for each pruning ratio {30%, 50%, 70%, 80%, 90%} we use different batch sizes with {128, 128, 128, 64, 32}. We set the initial learning rate as 0.1, the optimizer as SGD with momentum 0.9, and the scheduler as cosine annealing scheduler with weight decay 0.0005. For training ImageNet, we use ResNet-34 as the network architecture. For all coresets with different pruning rates, we train models for 300,000 iterations with a 256 batch size. We use the SGD optimizer with 0.9 momentum and 0.0001 weight decay, using a 0.1 initial learning rate. The cosine annealing learning rate scheduler was used for training. For a fair comparison, we used the same parameters across all pruning methods, including ours. All experiments were conducted using an NVIDIA A6000 GPU. We also attach the implementation in the supplementary material.
For calculating the DUAL score, we need three parameters $T$, $J$, and $c_{\mathcal{D}}$, each means score computation epoch, the length of the sliding window, and hyperparameter regarding the training dataset. We fix $J$ as 10 for all experiments. We use ($T$, $J$, $c_{\mathcal{D}}$) for each dataset as follows. For CIFAR-10, we use (30, 10, 5.5), for CIFAR-100, (30, 10, 4), and for ImageNet-1k, (60, 10, 11). We first roughly assign the term $c_{\mathcal{D}}$ based on the size of the initial dataset and by considering the relative difficulty of each, we set $c_{\mathcal{D}}$ for CIFAR-100 smaller than that of CIFAR-10. For the ImageNet-1k dataset, which contains 1,281,167 images, the size of the initial dataset is large enough that we do not need to set $c_{\mathcal{D}}$ to a small value to intentionally sample easier samples. Also, note that we fix the value of $C$ of Beta distribution at 15 across all experiments. A more detailed distribution, along with visualization, can be found in Appendix C.
Experiments with label noise and image corruption on CIFAR-100 are conducted under the same settings as described above, except for the hyperparameters for DUAL pruning. For label noise experiments, we set $T$ to 50 and $J$ to 10 across all label noise ratios. For $c_{\mathcal{D}}$, we set it to 6 for 20% and 30% noise, 8 for 40% noise. For image corruption experiments, we set $T$ to 30, $J$ to 10, and $c_{\mathcal{D}}$ to 6 across all image corruption ratios.
@@ Appendix B More Results on Experiments
We evaluate our proposed DUAL score through a wide range of analyses in this section. In Appendix B.1 and B.2, we demonstrate the robustness of the DUAL score across a variety of experiments. In Appendix B.3, we investigate the cross-architecture performance of DUAL pruning. In Appendix B.4, we show that beta sampling performs well even when combined with other score metrics, such as EL2N and Dyn-Unc, and also shows strong performance when compared to other sampling methods, especially CCS.
We first investigate the stability of our DUAL score compared to other baselines. We calculate the Spearman rank correlation of each score and the average score across five runs, following Paul et al. (2021). As shown in Figure 7, snapshot-based methods such as EL2N and Entropy exhibit relatively low correlation compared to methods that consider training dynamics. In particular, the DUAL score shows minimal score variation across runs, resulting in a high Spearman rank correlation, indicating strong stability across random seeds. Notably, even when the scores are calculated at the 30th epoch, the Spearman correlation between the individual scores and the overall average score remains approximately 0.95.
Figure 7: Average of Spearman rank correlation among independent runs and an overall average of five runs.
Next, we compute the Dyn-Unc, TDDS, and AUM scores at the 30th epoch, as we do for our method, and then compare the test accuracy on the coreset. Our pruning method, using the DUAL score and ratio-adaptive beta sampling, outperforms the others by a significant margin, as illustrated in Figure 8. We see using epoch 30 results in insufficient training dynamics for the others, thus it negatively impacts their performance.
@@
B.1 Image Classification with Label Noise
@@
We evaluated the performance of our proposed method across a wide range of pruning levels, from 10% to 90%, and compared the final accuracy with that of baseline methods. As shown in the Table 5-8, our method consistently outperforms the competition with a substantial margin in most cases. For a comprehensive analysis of performance under noisy conditions, please refer to Tables 5 to 7 for CIFAR-100, which show results for 20%, 30%, and 40% noise, respectively. Additionally, the results for 20% label noise in Tiny-ImageNet are in Table 8.
@@
DUAL+$\beta$ 45.21 ±$\pm 0.08$ 47.76 ±$\pm 0.33$ 48.99 ±$\pm 0.32$ 46.95 ±$\pm 0.23$ 43.01 ±$\pm 0.43$ 37.91 ±$\pm 0.28$ 28.78 ±$\pm 0.57$
B.2 Image Classification with Image Corruption
We also evaluated the robustness of our proposed method against five different types of realistic image corruption: motion blur, fog, reduced resolution, rectangular occlusion, and Gaussian noise across the corruption rate from 20% to 40%. The ratio of each type of corruption is 4% for 20% corruption, 6% for 30% corruption, and 8% for 40% corruption. Example images for each type of corruption can be found in Figure 12. Motion blur, reduced resolution, and rectangular occlusion are somewhat distinguishable, whereas fog and Gaussian noise are difficult for the human eye to differentiate. Surprisingly, our DUAL pruning prioritizes removing the most challenging examples, such as fog and Gaussian corrupted images, as shown in Figure 14.
@@ B.2 Image Classification with Image Corruption
We evaluated the performance of our proposed method across a wide range of pruning levels, from 10% to 90%, and compared the final accuracy with that of baseline methods. As shown in the table, our method consistently outperforms the competitors in most cases. For a comprehensive analysis of performance under noisy conditions, please refer to Tables 9 to 11 for CIFAR-100, which show results for 20%, 30%, and 40% corrupted images, respectively. Additionally, the results for 20% image corruption in Tiny-ImageNet are shown in Table 12.
@@ B.3 Cross-architecture generalization
In this section, we investigate the cross-architecture generalization ability of our proposed method. We calculate the example score on one architecture and test its coreset performance on a different architecture. This evaluation aims to assess the ability of example scores to be transferred across diverse architectural designs.
@@
Table 17: Cross-architecture generalization performance on CIFAR-100 from ResNet-18 to ResNet-50. We report an average of five runs. ‘R50 $\rightarrow$ R50’ stands for score computation on ResNet-50, as a baseline.
ResNet-18 $\rightarrow$ ResNet-50
Pruning Rate ($\rightarrow$) 30% 50% 70% 90%
Random 74.47 ±$\pm 0.67$ 70.09 ±$\pm 0.42$ 60.06 ±$\pm 0.99$ 41.91 ±$\pm 4.32$
EL2N 76.42 ±$\pm 1.00$ 69.14 ±$\pm 1.00$ 45.16 ±$\pm 3.21$ 19.63 ±$\pm 1.15$
Dyn-Unc 77.31 ±$\pm 0.34$ 72.12 ±$\pm 0.68$ 59.38 ±$\pm 2.35$ 31.74 ±$\pm 2.31$
CCS 74.78 ±$\pm 0.66$ 69.98 ±$\pm 1.18$ 59.75 ±$\pm 1.41$ 41.54 ±$\pm 3.94$
DUAL 78.03 ±$\pm 0.83$ 72.82 ±$\pm 1.46$ 63.08 ±$\pm 2.45$ 33.65 ±$\pm 2.92$
DUAL +$\beta$ sampling 77.82 ±$\pm 0.65$ 73.98 ±$\pm 0.62$ 66.36 ±$\pm 1.66$ 49.90 ±$\pm 2.56$
[…]
@@
B.4 Effectiveness of Beta Sampling

### v3

@@ A.2 Detailed Experimental Settings
Here we clarify the technical details in our works. For training the model on full-dataset and the selected subset, all parameters are used identically only except for batch sizes. For CIFAR-10/100, we train ResNet-18 for 200 epochs with batch size of 128, for each pruning ratio {30%, 50%, 70%, 80%, 90%} we use different batch sizes with {128, 128, 128, 64, 32}. We set the initial learning rate as 0.1, optimizer as SGD with momentum 0.9, and scheduler as cosine annealing scheduler with weight decay 0.0005. For training ImageNet, we use ResNet-34 as the network architecture. For all coresets with different pruning rates, we train models for 300,000 iterations with a 256 batch size. We use the SGD optimizer with 0.9 momentum and 0.0001 weight decay, using a 0.1 initial learning rate. The cosine annealing learning rate scheduler was used for training. For fair comparison, we use the same parameters across all pruning methods, including ours. All experiments were conducted using an NVIDIA A6000 GPU. We also attach the implementation in the supplementary material.
For calculating DUAL score, we need three parameters $T$, $J$, and $c_{\mathcal{D}}$, each means score computation epoch, the length of sliding window, and hyperparameter regarding the train dataset. We fix $J$ as 10 for all experiments. We use ($T$, $J$, $c_{\mathcal{D}}$) for each dataset as followings. For CIFAR-10, we use (30, 10, 5.5), for CIFAR-100, (30, 10, 4), and for ImageNet-1K, (60, 10, 11). We first roughly assign the term $c_{\mathcal{D}}$ based on the size of initial dataset and by considering the relative difficulty of each, we set $c_{\mathcal{D}}$ for CIFAR-100 smaller than that of CIFAR-10. For the ImageNet-1K dataset, which contains 1,281,167 images, the size of the initial dataset is large enough that we do not need to set $c_{\mathcal{D}}$ to a small value to in order to intentionally sample easier samples. Also, note that we fix the value of $C$ of Beta distribution at 15 across all experiments. A more detailed distribution, along with a visualization, can be found in Appendix C.
Experiments with label noise and image corruption on CIFAR-100 are conducted under the same settings as described above, except for the hyperparameters for DUAL pruning. For label noise experiments, we set $T$ to 50 and $J$ to 10 across all label noise ratio. For $c_{\mathcal{D}}$, we set it to 6 for 20% and 30% noise, 8 for 40% noise. For image corruption experiments, we set $T$ to 30, $J$ to 10, and $c_{\mathcal{D}}$ to 6 across all image corruption ratio.
@@ Appendix B More Results on Experiments
We evaluate our proposed DUAL score through a wide range of analyses in this section. In Appendix B.1 and B.2, we demonstrate the robustness of the DUAL score through intensive experiments. In Appendix B.3, we investigate the cross-architecture performance of our method. In Appendix B.6, we show the effectiveness of our Beta sampling when combined with other existing scores, compared to previous sampling strategies.
We first investigate the stability of our DUAL score. We calculate the Spearman rank correlation of individual scores and the average across five runs, following Paul et al. (2021). As shown in Figure 7, snapshot-based methods such as EL2N and Entropy exhibit relatively low correlation compared to methods using training dynamics. In particular, the DUAL score shows minimal variation across runs with a high Spearman rank correlation. This shows strong stability across random seeds.
Figure 7: Average of Spearman rank correlation among independent runs and overall average of five runs. DUAL score is calculated at 30th epoch.
Next, we compute the Dyn-Unc, TDDS, and AUM scores at the 30th epoch, as we do for our method, and then compare the test accuracy on the coreset. Our pruning method, using the DUAL score and ratio-adaptive Beta sampling, outperforms the others by a significant margin, as illustrated in Figure 8. We see that using epoch of 30 results in insufficient training dynamics for the others, thus it negatively impacts their performance.
@@
Figure 10 is a visualization of samples kept and pruned by our method. Samples kept by our method are more recognizable. The black swan on the grass and the sun-shaped balloon are rare cases, while the others are more easily recognizable. Samples pruned by our method are either typical, confusing, or mislabeled. The first and fifth examples are mislabeled, the second and third are typical, and the fourth is confusing.
(a) Samples kept by our method.
(b) Samples pruned by our method.
Figure 10: Illustrations of samples kept and pruned by our method at the pruning ratio of 30%. The pruned samples are likely either typical, confusing, or mislabeled, while the kept ones are certainly recognizable.
We compare the subset selected at the high pruning ratios by previous SOTA methods, namely CCS and D2. First, we examine the total amount of overlap by counting the number of samples in the intersection of each method in Table 5. Furthermore, for intuitive understanding, we visualize the selected subset by each method for the 70%, 80% pruning cases on CIFAR-100 in Figure 11. We can see that DUAL+$\beta$ seems to include more difficult examples than others.
Table 5: Comparison of subset overlap ratio over different pruning methods at high pruning rates. Let $S$ and $T$ denote the subsets selected by the two methods. The subset overlap ratio is computed as $\frac{|S\cap T|}{|S|}$ (which is equal to $\frac{|S\cap T|}{|T|}$).
Pruning Rate CCS & DUAL+$\beta$ D2 & DUAL+$\beta$ CCS & D2
70% 0.41 0.37 0.42
80% 0.31 0.27 0.34
90% 0.19 0.17 0.21
Figure 11: Comparison of selected subset over different methods at high pruning rates.
B.1 Image Classification Under Label Noise
@@
We evaluated the performance of our proposed method across a wide range of pruning levels, from 10% to 90%, and compared the final accuracy with that of baseline methods. As shown in the Table 6-9, our method consistently outperforms the competition with a substantial margin in most cases. For a comprehensive analysis of performance under noisy conditions, please refer to Tables 6 to 8 for CIFAR-100, which show results for 20%, 30%, and 40% noise, respectively. Additionally, the results for 20% label noise in Tiny-ImageNet are shown in Table 9.
@@
DUAL+$\beta$ sampling 45.21 ±$\pm 0.08$ 47.76 ±$\pm 0.33$ 48.99 ±$\pm 0.32$ 46.95 ±$\pm 0.23$ 43.01 ±$\pm 0.43$ 37.91 ±$\pm 0.28$ 28.78 ±$\pm 0.57$
B.2 Image Classification Under Image Corruption
We also evaluated the robustness of our proposed method against five different types of realistic image corruption: motion blur, fog, reduced resolution, rectangular occlusion, and Gaussian noise across the corruption rate from 20% to 40%. The ratio of each type of corruption is 4% for 20% corruption, 6% for 30% corruption, and 8% for 40% corruption. Example images for each type of corruption can be found in Figure 14. Motion blur, reduced resolution, and rectangular occlusion are somewhat distinguishable, whereas fog and Gaussian noise are difficult for the human eye to differentiate. Somewhat surprisingly, our DUAL pruning prioritize to remove the most challenging examples, such as fog and Gaussian corrupted images, as shown in Figure 16.
@@ B.2 Image Classification Under Image Corruption
We evaluated the performance of our proposed method across a wide range of pruning levels, from 10% to 90%, and compared the final accuracy with that of baseline methods. As shown in the table, our method consistently outperforms the competitors in most cases. For a comprehensive analysis of performance under noisy conditions, please refer to Tables 10 to 12 for CIFAR-100, which show results for 20%, 30%, and 40% corrupted images, respectively. Additionally, the results for 20% image corruption in Tiny-ImageNet are shown in Table 13.
@@ B.3 Cross-architecture generalization
In this section, we investigate the cross-architecture generalization ability of our proposed method. Specifically, we calculate the example score on one architecture and test its coreset performance on a different architecture. This evaluation, with results presented in Tables 14 through 18, aims to assess the transferability of these scores across diverse architectural designs.
@@
B.4 Image Classification on Long-tailed Distributions
We also conducted experiments on long-tailed versions of CIFAR-10 and CIFAR-100, following the procedure of Cao et al. (2019). These long-tailed datasets were constructed using an imbalance ratio, $\rho$, defined as the ratio between the sample sizes of the most frequent class ($n_{\max}$) and the least frequent class ($n_{\min}$), i.e. $\rho=n_{\max}/n_{\min}$. The class distribution in this long-tailed setup exhibits an exponential decay in sample sizes across classes. For all pruning ratios, we compared our method against several baselines: Random, EL2N, Dyn-Unc, and CCS. The results presented in Table 19 and 20 demonstrate that DUAL pruning (along with the Beta sampling) achieves superior performance compared to other baselines.
Table 19: Test accuracy on long-tailed imbalance on CIFAR-10. The test accuracy on a full dataset is 89.98 $(\rho=10)$ and 75.03 $(\rho=100)$. We report the average performance across three runs.
CIFAR-10-LT
Imbalance Ratio 10 100
Pruning Rate 30% 50% 70% 80% 90% 30% 50% 70% 80% 90%
Random 42.48 ±$\pm 0.45$ 28.20 ±$\pm 0.07$ 18.85 ±$\pm 0.24$ 10.00 ±$\pm 0.00$ 10.00 ±$\pm 0.00$ 28.23 ±$\pm 0.09$ 19.36 ±$\pm 0.18$ 10.00 ±$\pm 0.00$ 10.00 ±$\pm 0.00$ 10.00 ±$\pm 0.00$
EL2N 89.42 ±$\pm 0.20$ 87.59 ±$\pm 0.97$ 68.15 ±$\pm 3.44$ 52.90 ±$\pm 1.87$ 33.25 ±$\pm 0.41$ 72.70 ±$\pm 1.58$ 66.06 ±$\pm 4.27$ 52.90 ±$\pm 2.88$ 41.79 ±$\pm 2.61$ 30.30 ±$\pm 0.58$
Dyn-Unc 89.64 ±$\pm 0.28$ 87.60 ±$\pm 0.39$ 67.60 ±$\pm 4.34$ 53.05 ±$\pm 0.88$ 39.16 ±$\pm 1.94$ 74.40 ±$\pm 1.32$ 70.22 ±$\pm 1.60$ 51.89 ±$\pm 3.08$ 41.27 ±$\pm 2.34$ 31.24 ±$\pm 0.23$
CCS 84.42 ±$\pm 0.89$ 73.04 ±$\pm 1.20$ 47.07 ±$\pm 0.68$ 37.38 ±$\pm 0.36$ 27.91 ±$\pm 0.96$ 63.18 ±$\pm 1.56$ 45.46 ±$\pm 1.33$ 32.66 ±$\pm 0.63$ 29.38 ±$\pm 0.71$ 24.10 ±$\pm 0.97$
DUAL 89.67 ±$\pm 0.40$ 88.75 ±$\pm 0.36$ 75.38 ±$\pm 3.41$ 56.70 ±$\pm 2.83$ 43.58 ±$\pm 2.45$ 72.94 ±$\pm 1.14$ 69.66 ±$\pm 0.73$ 52.80 ±$\pm 1.00$ 38.32 ±$\pm 1.28$ 25.30 ±$\pm 1.28$
DUAL + $\beta$ 89.49 ±$\pm 0.21$ 88.12 ±$\pm 0.61$ 76.00 ±$\pm 2.79$ 78.31 ±$\pm 2.26$ 71.27 ±$\pm 1.44$ 73.81 ±$\pm 2.06$ 68.89 ±$\pm 0.24$ 52.95 ±$\pm 2.79$ 46.49 ±$\pm 1.80$ 36.43 ±$\pm 1.00$
Table 20: Test accuracy on long-tailed imbalance on CIFAR-100. The test accuracy on a full dataset is 62.92 $(\rho=10)$ and 41.67 $(\rho=100)$. We report the average performance across three runs.
CIFAR-100-LT
Imbalance Ratio 10 100
Pruning Rate 30% 50% 70% 80% 90% 30% 50% 70% 80% 90%
Random 32.89 ±$\pm 0.23$ 18.79 ±$\pm 0.75$ 8.26 ±$\pm 0.41$ 5.43 ±$\pm 0.07$ 3.23 ±$\pm 0.05$ 22.88 ±$\pm 0.87$ 11.45 ±$\pm 0.11$ 5.90 ±$\pm 0.15$ 3.96 ±$\pm 0.05$ 2.48 ±$\pm 0.02$
EL2N 57.57 ±$\pm 0.50$ 47.23 ±$\pm 0.46$ 21.38 ±$\pm 0.33$ 13.92 ±$\pm 0.97$ 9.54 ±$\pm 0.19$ 37.59 ±$\pm 2.13$ 24.76 ±$\pm 1.87$ 12.33 ±$\pm 0.54$ 9.42 ±$\pm 0.26$ 6.64 ±$\pm 0.02$
Dyn-Unc 58.09 ±$\pm 0.85$ 46.68 ±$\pm 0.69$ 25.95 ±$\pm 2.16$ 20.80 ±$\pm 0.68$ 13.48 ±$\pm 0.50$ 37.82 ±$\pm 1.08$ 26.88 ±$\pm 0.38$ 15.41 ±$\pm 0.42$ 12.47 ±$\pm 0.55$ 9.52 ±$\pm 0.12$
CCS 46.51 ±$\pm 0.56$ 34.85 ±$\pm 0.79$ 18.08 ±$\pm 0.80$ 11.34 ±$\pm 0.30$ 6.06 ±$\pm 0.43$ 27.46 ±$\pm 0.27$ 17.85 ±$\pm 0.76$ 11.43 ±$\pm 0.33$ 8.25 ±$\pm 0.66$ 4.34 ±$\pm 0.53$
DUAL 58.50 ±$\pm 0.27$ 54.11 ±$\pm 0.27$ 39.15 ±$\pm 1.43$ 30.10 ±$\pm 0.97$ 18.80 ±$\pm 1.17$ 36.35 ±$\pm 0.66$ 30.19 ±$\pm 1.58$ 20.47 ±$\pm 0.30$ 17.76 ±$\pm 0.47$ 12.52 ±$\pm 0.56$
DUAL + $\beta$ 58.05 ±$\pm 0.34$ 54.88 ±$\pm 0.36$ 43.53 ±$\pm 0.66$ 35.87 ±$\pm 1.75$ 27.13 ±$\pm 1.49$ 37.04 ±$\pm 0.97$ 32.25 ±$\pm 0.45$ 21.94 ±$\pm 1.27$ 19.38 ±$\pm 0.77$ 15.42 ±$\pm 0.32$
B.5 Comparison with Dynamic Pruning Methods
In this section, we present several experiments comparing recent dynamic pruning methods, such as those by Yuan et al. (2025) and Qin et al. (2024), with static approaches, including DUAL pruning. We first highlight two key differences between static and dynamic data pruning.
  • Compared to static pruning, dynamic pruning maintains access to the entire original dataset throughout training, allowing it to fully leverage all available information in the original dataset.
  • While both aim to improve training efficiency, their underlying goals differ slightly. Static data pruning seeks to identify a “fixed” subset that reduces the dataset size while preserving as much information about the original dataset as possible. This subset can then serve as a new, independent dataset, reusable across various model architectures and experimental setups. In contrast, dynamic data pruning enhances training efficiency within a single training session by pruning data dynamically on the fly. However, this approach requires storing the entire original dataset, making dynamic pruning less memory-efficient and not reusable.
Standard Training.
We conducted experiments on CIFAR-10 and CIFAR-100 with ResNet-18, using the same hyperparameters as described in Section 4. We first tested dynamic random pruning, which dynamically prunes randomly selected samples from the entire dataset at each epoch. Notably, dynamic random pruning significantly outperformed all static baselines, achieving test accuracies of 91.82% on CIFAR-10 and 72.8% on CIFAR-100 at a pruning ratio of 90%. We also evaluated the methods in Yuan et al. (2025); Qin et al. (2024) and the results are provided in Figure 17. Overall, dynamic methods consistently outperform static baselines. However, at lower pruning ratios (e.g., on CIFAR-10), DUAL can outperform dynamic methods under a similar computational budget.
We believe this performance gap stems from differences in accessible information: static methods are limited to 10% of the data, while dynamic methods use the full dataset. Consequently, the performance gap widens even further at aggressive pruning ratios. To validate this, we plot how often each sample was seen during training. The plot in Figure 18 shows that static methods are confined to a subset, while dynamic ones use nearly all data—rendering direct comparison somewhat unfair. Indeed, dynamic pruning methods might be better compared with scheduled batch-selection approaches, such as curriculum learning, rather than static pruning methods.
Label Noise Setting.
We also evaluated these methods under label noise conditions. In fact, Yuan et al. (2025) conclude that IES cannot prune any samples (corrupted or not) when label noise is introduced. Similarly, InfoBatch (Qin et al., 2024) tends to retain harder (and often noisy) samples, as it removes only easy examples during training. In contrast, DUAL effectively filters noisy samples, improving performance even beyond full-data training.
We conducted experiments on CIFAR-100 with a 40% label noise setting (full-train test accuracy: 52.74%) to verify this explanation. DUAL achieves over 70% test accuracy at a 50% pruning ratio as can be seen in Table 8 in Appendix B.1, whereas InfoBatch achieves only 51.24% accuracy with a similar number of iterations. Under similar iterations, random dynamic pruning achieves 51.81% test accuracy, which still outperforms random static pruning. Lastly, IES (Yuan et al., 2025) prunes only 1.7% of samples during training (consistent with the original report in their paper), resulting in 51.95% test accuracy. Furthermore, our static method can create fixed subsets in which nearly all noisy samples have been removed, resulting in high-quality datasets that can be preserved for future use.
Figure 17: This figure shows test accuracy (y-axis) versus the total number of iterations needed to fully train each subset (x-axis). The top figure corresponds to CIFAR-10, and the bottom to CIFAR-100. Results are averaged over five random seeds. For the static pruning methods, we plotted the number of iterations and test accuracy at pruning ratios of 30%, 50%, 70%, 80%, and 90%, which correspond to the markers on each line from right to left. We adjust the total training epochs of InfoBatch following the procedure in the original paper, in order to evaluate its performance when the total number of iterations is reduced. For IES, we adjust the pruning threshold to evaluate its performance under a reduced total number of training iterations. Results demonstrate that dynamic pruning methods—including even random dynamic pruning—outperform static baselines in accuracy while requiring fewer iterations. Among static methods, DUAL is both the most efficient and the best-performing. However, it does not achieve as favorable a time-performance trade-off as dynamic methods. Still, at low pruning ratios—where more information is available—DUAL performs comparably to dynamic approaches.
Figure 18: This figure shows the selection frequency of each sample at a high pruning ratio (90%). The x-axis represents data indices sorted in descending order by selection frequency, and the y-axis indicates the normalized selection ratio—i.e., how often each data point was selected, divided by the total number of data points seen during training (values sum to 1). Methods like Random Dynamic Pruning, InfoBatch, and IES use nearly the entire dataset during training, making them incomparable to static pruning methods, which are restricted to only 10% of the data.
B.6 Ablation Study on Beta Sampling